OpenStudy (anonymous):

The sped of a passenger train is 14mph faster than the speed of a freight train. The passenger train travels 280 miles in the same time it takes the freight train to travel 210 miles. Find the speed of each train.

OpenStudy (anonymous):

In these written questions the hard part is to find the equations. Let x be the speed of the slower train, than the faster trains speed is x+14. Now the distance travelled is speed*time. So we have two equations (t is time) x*t=210 (x+14)t=280 dividing the two equations give (x+14)/x=280/210 Hope you can solve this.